An organic compound contains 88.89% carbon and 11.11% hydrogen by mass. The molecular mass of the organic compounds is 54g:mol. Find the molecular formula of the compound.
1
Expert's answer
2020-06-08T15:24:40-0400

Solution.

"w(C)=88.89\%;"

"w(H)=11.11\%;"

"M(C_xH_y)=54g\/mol;"

"x:y=\dfrac{w(C)}{Ar(C)}:\dfrac{w(H)}{Ar(H)}=\dfrac{88.89}{12}:\dfrac{11.11}{1}=7.4075:11.11=1:1.5=10:15=2:3;"

"C_2H_3;"

"M(C_2H_3)=2\sdot12+3=27g/mol;"

Since "M(C_2H_3)=27g\/mol" , and in the condition 54g / mol, so the indices need to be doubled.

"C_4H_6;"

"M(C_4H_6)=12\sdot4+6=54g/mol;"

Answer: "C_4H_6."